OpenStudy (hba):

The point (3,2) lies on the graph of the inverse of f(x)=2x^3+x+A.The Value of A is :

OpenStudy (hba):

@hartnn

hartnn (hartnn):

then (2,3) will lie on that graph

OpenStudy (hba):

f(x)=0

OpenStudy (hba):

So -18 is the answer ? @hartnn

hartnn (hartnn):

3=18+A

OpenStudy (hba):

^ please Explain ? @hartnn

hartnn (hartnn):

put 2,3 in that equation

OpenStudy (hba):

f(x)=2(2)^3*2+A

OpenStudy (hba):

f(x)=18+A

OpenStudy (hba):

What next ? @hartnn

hartnn (hartnn):

3= 18+A

hartnn (hartnn):

f(x) =3 when x = 2

OpenStudy (hba):

So how come f(x)=3 ?

OpenStudy (hba):

@hartnn

hartnn (hartnn):

to put point (2,3) means f(x) =3 when x = 2

OpenStudy (hba):

Thanks Alot.